OpenStudy (anonymous):

can someone please explain these to me? My book makes this sooo complicated! f(x)= 2x^3-5x; g(x)=x^2+8x+15 what is the first step?

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

Find f + g and its domain.

OpenStudy (ranga):

Just add f(x) and g(x)

OpenStudy (anonymous):

oh! 2x^3 + x ^2+ 3x + 15

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the domain of the combination has to be evaluated by itself. there are norestriction for values which x can be, so the domain of (f+g)(x) = (-infinity, infinity)

OpenStudy (anonymous):

yay are there any restrictions or how can you tell if there is a restriction?

OpenStudy (ranga):

Domain is all x values that are allowed. If a value of x would make f(x) undefined then that x value should be excluded. For example, f(x) = 1 / (x - 2) is not defined when x = 2. So the domain should exclude x = 2.

OpenStudy (anonymous):

got it =) thank you

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

if the next thing they ask me for is f-g do I just switch the positive signs to negative?